OpenStudy (anonymous):

Solve the following formula for x FOR MEDAL PLZZZ HELP y = 4 mx -7 b x=

OpenStudy (anonymous):

Are you sure you are solving for the variable x, or the slope?

OpenStudy (anonymous):

in the question its telling me formula

OpenStudy (anonymous):

variable x

OpenStudy (anonymous):

Is the equation y=4x-7?

OpenStudy (anonymous):

y = 4x-7 4x-7 = y +7 +7 4x = y+7 divide both sides by 4 x = (y+7)/4

OpenStudy (anonymous):

thankx

OpenStudy (anonymous):

Solve the following formula for x y = 8 mx + 6 b x=

OpenStudy (anonymous):

how about this one